suppose he has a standard normal distribution with a mean of zero and standard deviation of one the probability that Z values are larger than blank is .3483
Answer #1

X ~ N ( µ = 0 , σ = 1 )
P ( X > x ) = 1 - P ( X < x ) = 1 - 0.3483 = 0.6517
To find the value of x
Looking for the probability 0.6517 in standard normal table to calculate Z score = 0.39
Z = ( X - µ ) / σ
0.3899 = ( X - 0 ) / 1
X = 0.39
P ( X > 0.39 ) = 0.3483